# Neutrino mixing parameters and masses from $\Delta(96)\rtimes H_{CP}$ in the tri-direct CP approach

### v1

@@ Appendix A Group theory of $\Delta(96)$
$\begin{array}[]{l}\bm{3_{m}}:\quad~{}~{}\begin{pmatrix}\alpha_{3}\beta_{1}+\alpha_{2}\beta_{6}\\ \alpha_{1}\beta_{2}+\alpha_{3}\beta_{5}\\ \alpha_{2}\beta_{3}+\alpha_{1}\beta_{4}\end{pmatrix}\\ \bm{3_{[m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{6}-\alpha_{3}\beta_{1}\\ \alpha_{3}\beta_{5}-\alpha_{1}\beta_{2}\\ \alpha_{1}\beta_{4}-\alpha_{2}\beta_{3}\end{pmatrix}\\ \bm{\hat{3}_{m}}:~{}\quad~{}\begin{pmatrix}\alpha_{3}\beta_{2}+\alpha_{2}\beta_{4}\\ \alpha_{1}\beta_{3}+\alpha_{3}\beta_{6}\\ \alpha_{2}\beta_{1}+\alpha_{1}\beta_{5}\end{pmatrix}\\ \bm{\hat{3}_{[m+1]}}:~{}\begin{pmatrix}\alpha_{2}\beta_{4}-\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{6}-\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{5}-\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{6}:P_{6}^{m}\begin{pmatrix}\alpha_{3}\beta_{3}\\ \alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ -\alpha_{1}\beta_{6}\\ -\alpha_{3}\beta_{4}\\ -\alpha_{2}\beta_{5}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\bar{3}_{m}}:~{}\quad~{}\begin{pmatrix}\alpha_{2}\beta_{2}+\alpha_{3}\beta_{4}\\ \alpha_{3}\beta_{3}+\alpha_{1}\beta_{6}\\ \alpha_{1}\beta_{1}+\alpha_{2}\beta_{5}\end{pmatrix}\\ \bm{\bar{3}_{[m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{2}-\alpha_{3}\beta_{4}\\ \alpha_{3}\beta_{3}-\alpha_{1}\beta_{6}\\ \alpha_{1}\beta_{1}-\alpha_{2}\beta_{5}\end{pmatrix}\\ \bm{\hat{3}_{m}}:~{}\quad~{}\begin{pmatrix}\alpha_{2}\beta_{1}+\alpha_{3}\beta_{6}\\ \alpha_{3}\beta_{2}+\alpha_{1}\beta_{5}\\ \alpha_{1}\beta_{3}+\alpha_{2}\beta_{4}\end{pmatrix}\\ \bm{\hat{3}_{[m+1]}}:~{}\begin{pmatrix}\alpha_{2}\beta_{1}-\alpha_{3}\beta_{6}\\ \alpha_{3}\beta_{2}-\alpha_{1}\beta_{5}\\ \alpha_{1}\beta_{3}-\alpha_{2}\beta_{4}\end{pmatrix}\\ \bm{6}:P_{6}^{m}\begin{pmatrix}\alpha_{1}\beta_{2}\\ \alpha_{2}\beta_{3}\\ \alpha_{3}\beta_{1}\\ -\alpha_{3}\beta_{5}\\ -\alpha_{2}\beta_{6}\\ -\alpha_{1}\beta_{4}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{3_{m}}:~{}\quad~{}\begin{pmatrix}\alpha_{3}\beta_{3}+\alpha_{2}\beta_{5}\\ \alpha_{1}\beta_{1}+\alpha_{3}\beta_{4}\\ \alpha_{2}\beta_{2}+\alpha_{1}\beta_{6}\end{pmatrix}\\ \bm{3_{[m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{5}-\alpha_{3}\beta_{3}\\ \alpha_{3}\beta_{4}-\alpha_{1}\beta_{1}\\ \alpha_{1}\beta_{6}-\alpha_{2}\beta_{2}\end{pmatrix}\\ \bm{\bar{3}_{m}}:~{}\quad~{}\begin{pmatrix}\alpha_{2}\beta_{3}+\alpha_{3}\beta_{5}\\ \alpha_{3}\beta_{1}+\alpha_{1}\beta_{4}\\ \alpha_{1}\beta_{2}+\alpha_{2}\beta_{6}\end{pmatrix}\\ \bm{\bar{3}_{[m+1]}}:~{}\begin{pmatrix}\alpha_{2}\beta_{3}-\alpha_{3}\beta_{5}\\ \alpha_{3}\beta_{1}-\alpha_{1}\beta_{4}\\ \alpha_{1}\beta_{2}-\alpha_{2}\beta_{6}\end{pmatrix}\\ \bm{6}:P_{2}^{m}\begin{pmatrix}\alpha_{2}\beta_{4}\\ \alpha_{3}\beta_{6}\\ \alpha_{1}\beta_{5}\\ -\alpha_{2}\beta_{1}\\ -\alpha_{1}\beta_{3}\\ -\alpha_{3}\beta_{2}\end{pmatrix}\end{array}$
@@
Table 8: Residual symmetry restrictions on flavon vacuum alignments for $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ transforming under either the $\bm{\bar{3}_{0}}$ or $\bm{\bar{3}_{1}}$. Both flavons develop null VEVs $(0,0,0)^{T}$ when preserving $K_{4}$, $Z_{8}$ and the three $Z_{4}$ subgroups $Z_{4}^{cd^{2}}$, $Z_{4}^{cd^{3}}$ and $Z_{4}^{c^{2}d^{3}}$. $\phi_{\text{sol}}\sim\bm{\bar{3}_{1}}$ similarly shows vanishing VEVs under the other nine $Z_{4}$ subgroups. Vacuum alignment for $\phi_{\text{sol}}\sim\bm{\bar{3}_{1}}$ under all $Z_{3}$ subgroups and the three $Z_{2}$ subgroups whose generators do not include generator $b$ match those in the $\bm{\bar{3}_{0}}$ case. While only one representative residual CP transformation is explicitly presented, all compatible CP-invariant VEV configurations can be generated through global phase factors $e^{\pm i\frac{\pi}{4}}$ or $i$ multiplication. These phase contributions become physically irrelevant through parameter redefinition $\eta\to\eta\pm\frac{\pi}{2}$ or $\eta+\pi$. Residual CP symmetry enforces real VEV magnitudes $v_{\phi_{a}}$ and $v_{\phi_{s}}$, with $x$ remaining a free real parameter.
@@
In the tri-direct CP approach, the original symmetry $\Delta(96)\rtimes H_{CP}$ breaks spontaneously into $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ (atmospheric sector) and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ (solar sector) through flavons $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$, where $G_{\text{atm}}$ and $G_{\text{sol}}$ are Abelian subgroups of $\Delta(96)$. These flavons adopt vacuum alignments are invariant under the action of the corresponding residual symmetries. Here, the flavon $\phi_{\text{atm}}$ transform as triplet representation $\bm{\bar{3}_{0}}$, and the flavon $\phi_{\text{sol}}$ is assigned to triplet representation $\bm{\bar{3}_{0}}$ or $\bm{\bar{3}_{1}}$. The most general VEVs of the flavons $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ which preserve the possible residual symmetries are summarized in table LABEL:tab:invariant_VEVs.
@@ Appendix B Diagonalization of the neutrino mass matrix $m^{\prime}_{\nu}$
$m^{\prime}_{\nu}=\begin{pmatrix}0&~{}0&~{}0\\ 0&~{}|y|e^{i\phi_{y}}~{}&~{}|z|e^{i\phi_{z}}\\ 0&~{}|z|e^{i\phi_{z}}~{}&~{}|w|e^{i\phi_{w}}\end{pmatrix}\,,$ (B.1)
@@
$U^{T}_{\nu 2}m^{\prime}_{\nu}U_{\nu 2}=\text{diag}(0,m_{2},m_{3})\,,$ (B.2)
@@
$m^{2}_{2}=\frac{1}{2}\left[|y|^{2}+|w|^{2}+2|z|^{2}-\frac{|w|^{2}-|y|^{2}}{\cos 2\theta}\right],\quad m^{2}_{3}=\frac{1}{2}\left[|y|^{2}+|w|^{2}+2|z|^{2}+\frac{|w|^{2}-|y|^{2}}{\cos 2\theta}\right]\,.$ (B.3)
The rotational parameter $\theta$, the phase parameters $\psi$, $\rho$ and $\sigma$ are determined by the following angular relations:

### v2

@@ Appendix A Group theory of $\Delta(96)$
$\begin{array}[]{l}\bm{3_{m}}:\qquad~~\begin{pmatrix}\alpha_{3}\beta_{1}+\alpha_{2}\beta_{6}\\ \alpha_{1}\beta_{2}+\alpha_{3}\beta_{5}\\ \alpha_{2}\beta_{3}+\alpha_{1}\beta_{4}\end{pmatrix}\\ \bm{3_{[m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{6}-\alpha_{3}\beta_{1}\\ \alpha_{3}\beta_{5}-\alpha_{1}\beta_{2}\\ \alpha_{1}\beta_{4}-\alpha_{2}\beta_{3}\end{pmatrix}\\ \bm{\hat{3}_{m}}:~\qquad~\begin{pmatrix}\alpha_{3}\beta_{2}+\alpha_{2}\beta_{4}\\ \alpha_{1}\beta_{3}+\alpha_{3}\beta_{6}\\ \alpha_{2}\beta_{1}+\alpha_{1}\beta_{5}\end{pmatrix}\\ \bm{\hat{3}_{[m+1]}}:~\begin{pmatrix}\alpha_{2}\beta_{4}-\alpha_{3}\beta_{2}\\ \alpha_{3}\beta_{6}-\alpha_{1}\beta_{3}\\ \alpha_{1}\beta_{5}-\alpha_{2}\beta_{1}\end{pmatrix}\\ \bm{6}:P_{6}^{m}\begin{pmatrix}\alpha_{3}\beta_{3}\\ \alpha_{1}\beta_{1}\\ \alpha_{2}\beta_{2}\\ -\alpha_{1}\beta_{6}\\ -\alpha_{3}\beta_{4}\\ -\alpha_{2}\beta_{5}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{\bar{3}_{m}}:~\qquad~\begin{pmatrix}\alpha_{2}\beta_{2}+\alpha_{3}\beta_{4}\\ \alpha_{3}\beta_{3}+\alpha_{1}\beta_{6}\\ \alpha_{1}\beta_{1}+\alpha_{2}\beta_{5}\end{pmatrix}\\ \bm{\bar{3}_{[m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{2}-\alpha_{3}\beta_{4}\\ \alpha_{3}\beta_{3}-\alpha_{1}\beta_{6}\\ \alpha_{1}\beta_{1}-\alpha_{2}\beta_{5}\end{pmatrix}\\ \bm{\hat{3}_{m}}:~\qquad~\begin{pmatrix}\alpha_{2}\beta_{1}+\alpha_{3}\beta_{6}\\ \alpha_{3}\beta_{2}+\alpha_{1}\beta_{5}\\ \alpha_{1}\beta_{3}+\alpha_{2}\beta_{4}\end{pmatrix}\\ \bm{\hat{3}_{[m+1]}}:~\begin{pmatrix}\alpha_{2}\beta_{1}-\alpha_{3}\beta_{6}\\ \alpha_{3}\beta_{2}-\alpha_{1}\beta_{5}\\ \alpha_{1}\beta_{3}-\alpha_{2}\beta_{4}\end{pmatrix}\\ \bm{6}:P_{6}^{m}\begin{pmatrix}\alpha_{1}\beta_{2}\\ \alpha_{2}\beta_{3}\\ \alpha_{3}\beta_{1}\\ -\alpha_{3}\beta_{5}\\ -\alpha_{2}\beta_{6}\\ -\alpha_{1}\beta_{4}\end{pmatrix}\end{array}$ $\begin{array}[]{l}\bm{3_{m}}:~\qquad~\begin{pmatrix}\alpha_{3}\beta_{3}+\alpha_{2}\beta_{5}\\ \alpha_{1}\beta_{1}+\alpha_{3}\beta_{4}\\ \alpha_{2}\beta_{2}+\alpha_{1}\beta_{6}\end{pmatrix}\\ \bm{3_{[m+1]}}:\begin{pmatrix}\alpha_{2}\beta_{5}-\alpha_{3}\beta_{3}\\ \alpha_{3}\beta_{4}-\alpha_{1}\beta_{1}\\ \alpha_{1}\beta_{6}-\alpha_{2}\beta_{2}\end{pmatrix}\\ \bm{\bar{3}_{m}}:~\qquad~\begin{pmatrix}\alpha_{2}\beta_{3}+\alpha_{3}\beta_{5}\\ \alpha_{3}\beta_{1}+\alpha_{1}\beta_{4}\\ \alpha_{1}\beta_{2}+\alpha_{2}\beta_{6}\end{pmatrix}\\ \bm{\bar{3}_{[m+1]}}:~\begin{pmatrix}\alpha_{2}\beta_{3}-\alpha_{3}\beta_{5}\\ \alpha_{3}\beta_{1}-\alpha_{1}\beta_{4}\\ \alpha_{1}\beta_{2}-\alpha_{2}\beta_{6}\end{pmatrix}\\ \bm{6}:P_{6}^{m}\begin{pmatrix}\alpha_{2}\beta_{4}\\ \alpha_{3}\beta_{6}\\ \alpha_{1}\beta_{5}\\ -\alpha_{2}\beta_{1}\\ -\alpha_{1}\beta_{3}\\ -\alpha_{3}\beta_{2}\end{pmatrix}\end{array}$
@@
Table 8: Residual symmetry restrictions on flavon vacuum alignments for $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ transforming under either the $\bm{\bar{3}_{0}}$ or $\bm{\bar{3}_{1}}$. Both flavons develop null VEVs $(0,0,0)^{T}$ when $K_{4}$, $Z_{8}$ and the three $Z_{4}$ subgroups $Z_{4}^{cd^{2}}$, $Z_{4}^{cd^{3}}$ and $Z_{4}^{c^{2}d^{3}}$ are preserved. $\phi_{\text{sol}}\sim\bm{\bar{3}_{1}}$ similarly shows vanishing VEVs under the other nine $Z_{4}$ subgroups. Vacuum alignment for $\phi_{\text{sol}}\sim\bm{\bar{3}_{1}}$ under all $Z_{3}$ subgroups and the three $Z_{2}$ subgroups whose generators do not include generator $b$ match those in the $\bm{\bar{3}_{0}}$ case. While only one representative residual CP transformation is explicitly presented, all compatible CP-invariant VEV configurations can be generated through global phase factors $e^{\pm i\frac{\pi}{4}}$ or $i$ multiplication. These phase contributions become physically irrelevant through parameter redefinition $\eta\to\eta\pm\frac{\pi}{2}$ or $\eta+\pi$. Residual CP symmetry enforces real VEV magnitudes $v_{\phi_{a}}$ and $v_{\phi_{s}}$, with $x$ remaining a free real parameter.
@@
In the tri-direct CP approach, the original symmetry $\Delta(96)\rtimes H_{CP}$ breaks spontaneously into $G_{\text{atm}}\rtimes H^{\text{atm}}_{CP}$ (atmospheric sector) and $G_{\text{sol}}\rtimes H^{\text{sol}}_{CP}$ (solar sector) through flavons $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$, where $G_{\text{atm}}$ and $G_{\text{sol}}$ are Abelian subgroups of $\Delta(96)$. These flavons adopt vacuum alignments which are invariant under the action of the corresponding residual symmetries. Here, the flavon $\phi_{\text{atm}}$ transforms as triplet representation $\bm{\bar{3}_{0}}$, and the flavon $\phi_{\text{sol}}$ is assigned to triplet representation $\bm{\bar{3}_{0}}$ or $\bm{\bar{3}_{1}}$. The most general VEVs of the flavons $\phi_{\text{atm}}$ and $\phi_{\text{sol}}$ which preserve the possible residual symmetries are summarized in table LABEL:tab:invariant_VEVs.
@@ Appendix B Diagonalization of the neutrino mass matrix $m^{\prime}_{\nu}$
$m^{\prime}_{\nu}=\left\{\begin{array}[]{ccc}\begin{pmatrix}0&~0&~0\\ 0&~|y|e^{i\phi_{y}}~&~|z|e^{i\phi_{z}}\\ 0&~|z|e^{i\phi_{z}}~&~|w|e^{i\phi_{w}}\end{pmatrix}&\mathrm{for}&\mathrm{NO\,,}\\ \\ \begin{pmatrix}|y|e^{i\phi_{y}}&|z|e^{i\phi_{z}}&~0\\ |z|e^{i\phi_{z}}&~|w|e^{i\phi_{w}}~&~0\\ 0&0~&~0\end{pmatrix}&\mathrm{for}&\mathrm{IO\,.}\end{array}\right$. (B.1)
@@
$U_{\nu 2}^{T}m_{\nu}^{\prime}U_{\nu 2}=\left\{\begin{array}[]{ccc}\mathrm{diag}(0,m_{2},m_{3})&\mathrm{for}&\mathrm{NO}\,,\\ \mathrm{diag}(m_{1},m_{2},0)&\mathrm{for}&\mathrm{IO}\,.\end{array}\right$. (B.2)
@@
$m^{2}_{i}=\frac{1}{2}\left[|y|^{2}+|w|^{2}+2|z|^{2}-\frac{|w|^{2}-|y|^{2}}{\cos 2\theta}\right],\quad m^{2}_{j}=\frac{1}{2}\left[|y|^{2}+|w|^{2}+2|z|^{2}+\frac{|w|^{2}-|y|^{2}}{\cos 2\theta}\right]\,.$ (B.3)
with $m_{i}$= $m_{2}$, $m_{j}$= $m_{3}$ for the NO case and $m_{i}$= $m_{1}$, $m_{j}$= $m_{2}$ for the IO case. The rotational parameter $\theta$, the phase parameters $\psi$, $\rho$ and $\sigma$ are determined by the following angular relations: